OpenStudy (anonymous):

if p+pq is 4 times p-pq, which of the following has exactly one value? (pq does not = 0) a)p b)q c)pq d)p+pq e)p-pq

OpenStudy (anonymous):

a) p

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

this is p+pq=4(p-pq) p+pq=4p-4pq pq+4pq=4p-p 5pq=3p this cancels p 5q=3 q=3/5

OpenStudy (anonymous):

so here p does not change

OpenStudy (anonymous):

if p+pq=2(p-pq) p+pq=2(p-pq) we can factor out p on both sides, gives p(1+q)=2p(1-q) p cancels here (1+q)=2(1-q) (1+q)=2-2q q+2q=2-1 3q=1 q=1/3 therfore q can change, p +q can change, and p-q also change, except p

OpenStudy (anonymous):

would that help you?

OpenStudy (anonymous):

yea thanks